OpenStudy (anonymous):

what is the greatest possible error for 11 1/8

OpenStudy (campbell_st):

isn't greatest possible error... 1/2 of the smallest unit of measurement... so your smallest unit is 1/8... so what's half of that..?

OpenStudy (anonymous):

like 1/4?

OpenStudy (campbell_st):

nope like \[\frac{1}{8} \times \frac{1}{2} = \]

OpenStudy (anonymous):

oh hahah 1/16

OpenStudy (anonymous):

thank ya :)